OpenStudy (anonymous):

good morning, can someone help me figure this out? Is this equation true or false? 28-4sqrt 2= 24sqrt 2

OpenStudy (anonymous):

false

OpenStudy (anonymous):

false

OpenStudy (anonymous):

Thank you, why is it false, Im having a hard time figuring out how to solve the equation

OpenStudy (anonymous):

your equation implies sqrt(2)=1 which is indeed false

OpenStudy (anonymous):

because the 4sqrt2 is connected, and therefore you would have to be divided to be separated.

OpenStudy (anonymous):

when you divide by 4 it would make the sides unequal

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

no worries

OpenStudy (anonymous):

just to make sure I fully understand, 28-4sqrt 2 ( the 4 and the two is connected) in order to make them seperated I would have to divide???? this is where I am confused...

OpenStudy (anonymous):

yes 4 and sqrt(2) is connected.

OpenStudy (anonymous):

I understand that much, its after that where I am confused

OpenStudy (anonymous):

divide by 4 if you want to separate rt(2) and 4

OpenStudy (anonymous):

to try to simplify this you would subtract the 28 from the left side to get: -4sqrt(2)=24sqrt(2)-28. [-4sqrt(2)]/-4=[24sqrt(2)-28]/4. so the final answer would be that sqrt(2) does not equal 6sqrt(2)-7

OpenStudy (anonymous):

Thank u....

OpenStudy (radar):

Try and solve it and you would discover: \[28-4\sqrt{2}=24\sqrt{2}\] \[28=24\sqrt{2}+4\sqrt{2}\] \[28=28\sqrt{2}\] Obviously that is false.